(3,8) and (9,12)
Determine in each case the gradient and equation of the line joining the two points.
Determine the co-ordinates for X intercept and y intercept for each of the lines

Answer by mananth(16946)

x1 y1 x2 y2
3 8 9 12

slope m =(y2-y1)/(x2-x1)
(12-8)/(9-3)
(4/6)
m=0.67

Plug value of the slope and point(3,8 ) in
Y = m x + b
8.00 = 2 + b
b= 8 - 2
b= 6
So the equation will be
Y =0.67 x+6
y intercept point (0,6)
x intercept point (8.95,0)